### v2

# How May U.S. Courts Scrutinize Their Recidivism Risk Assessment Tools? Contextualizing AI Fairness Criteria on a Judicial Scrutiny-based Framework

Abstract
The AI/HCI and legal communities have developed largely independent conceptualizations of fairness. This conceptual difference hinders the potential incorporation of technical fairness criteria (e.g., procedural, group, and individual fairness) into sustainable policies and designs, particularly for high-stakes applications like recidivism risk assessment. To foster common ground, we conduct legal research to identify if and how technical AI conceptualizations of fairness surface in primary legal sources. We find that while major technical fairness criteria can be linked to constitutional mandates such as “Due Process” and “Equal Protection” thanks to judicial interpretation, several challenges arise when operationalizing them into concrete statutes/regulations. These policies often adopt procedural and group fairness but ignore the major technical criterion of individual fairness. Regarding procedural fairness, judicial “scrutiny” categories are relevant but may not fully capture how courts scrutinize the use of demographic features in potentially discriminatory government tools like RRA. Furthermore, some policies contradict each other on whether to apply procedural fairness to certain demographic features. Thus, we propose a new framework, integrating U.S. demographics-related legal scrutiny concepts and technical fairness criteria, and contextualize it in three other major AI-adopting jurisdictions (EU, China, and India).
Introduction
Recidivism risk assessment (RRA) tools, widely deployed models that predict an individual’s likelihood to re-offend following a criminal charge, raise ethical concerns surrounding fairness, transparency, and potential bias. These tools rely on quantifiable factors (e.g., criminal history, age at first offense) to produce a risk score that assists court officials in bail, sentencing, parole, and other correctional decisions.
Outcomes of RRA systems have a direct impact on the lives of people accused or convicted of crime and an indirect impact on their families and communities: manifesting in the forms of emotional and economic hardship when incarcerated family members serve longer sentences or parole conditions, and in the forms of diminished social capital, disintegration, and cycles of crime and poverty for communities with high rates of incarceration.
RRAs have traditionally been studied in several fields, namely Psychology (Hanson and Morton-Bourgon 2009), Criminology (Caudy, Durso, and Taxman 2013), Law (Nishi 2019), Statistics (Imrey and Dawid 2015), and other interdisciplinary communities (Mann, Hanson, and Thornton 2010; Sreenivasan et al. 2000; Hamilton et al. 2016). In this cross-disciplinary work, spanning Law, AI and Human-Computer Interaction (HCI), we contribute new legal insights for fairness audit of RRA tools.
Research on the fairness of RRA tools has grown significantly across many disciplines since 2016, when an investigative article [^1] https://www.propublica.org/article/machine-bias-risk-assessments-in-criminal-sentencing from a non-profit organization, ProPublica, drew public attention to racial bias in COMPAS, a proprietary AI-based RRA tool used in Florida and many other U.S. states (Angwin et al. 2022).
Academic research on fairness gained much traction from multiple angles following that publicized project, sometimes by directly criticizing the use of AI for RRA on fairness grounds (Dressel and Farid 2018; Green and Chen 2019), other times by subtly pointing out nuanced limitations of fairness criteria (e.g., it is almost impossible to simultaneously optimize for several fairness criteria and/or accuracy (Chouldechova 2017; Berk et al. 2021)).
The technical AI literature includes procedural fairness (“fairness through unawareness” or excluding a feature from model input (Bart, Teodorescu, and Morse 2024)) and two mathematical categories of outcome-based fairness: group and individual fairness. Aligned with the intuitive definition by Kleinberg, Mullainathan, and Raghavan (2017), group fairness, or group parity, is achieved when a statistical metric of interest, e.g., positive outcome rate, is equalized across different groups with respect to a sensitive feature, e.g. race or sex/gender (Pedreshi, Ruggieri, and Turini 2008). Dwork et al. (2012) introduced individual fairness based on the intuition that similar individuals should get similar outcomes. Formally, for a pair of individuals, given an input-space distance metric to measure how differently situated they are and an output-space distance metric to measure how different the distributions of their possible outcomes are, individual fairness requires that their output-space distance should be upper-bound by their input-space distance. The more similarly situated two persons are (e.g., same demographics and criminal histories), the more likely that they receive similar outcomes (e.g., RRA scores).
RRA tools have already been widely adopted across U.S. government bodies for decades. Motivated by the emerging need to identify a constitutional basis for establishing enforceable fairness standards on such AI tools, our main research questions (RQs) are:
RQ1.a: Do technical AI fairness criteria for RRA have an enforceable legal basis under the U.S. Constitution?
RQ1.b: Are there any legal challenges when such abstract fairness criteria are translated into concrete policies?
RQ2: How to develop a new legal framework to mitigate those challenges, informed by established legal theories?
Two main contributions follow. First, we analyzed primary sources of law, including the U.S. Constitution, case law interpreting the Constitution, as well as statutes and regulations at the state and federal levels. The goal is to explore a legal basis for technical fairness criteria, as well as challenges when operationalizing legally relevant concepts of fairness. Second, to address the challenges, we propose an integrated framework that combines technical fairness criteria and the judicial “scrutiny” concept (typically applied when the government treats people differently based on their demographics). This “scrutiny” concept is particularly suited to the RRA use case (compared to other high-stakes AI-assisted decision-making use cases such as loan approval (Goyal et al. 2024) or job application screening (Rigotti and Fosch-Villaronga 2024), which are typically undertaken by private-sector actors) because the U.S. judiciary has developed elaborate scrutiny standards to scrutinize discriminations by public-sector actors (e.g., state and federal government bodies) against private citizens. Our framework utilizes a common unit of analysis, scrutiny-worthiness, to help auditors decide on which fairness-related criteria and for which demographics they should prioritize their audit.
Background: Conceptualization of Fairness
HCI vs. AI
For a high-level overview, Narayanan et al. (2024) conducted a comprehensive review of empirical research on the perceived fairness of AI-assisted decision-making in the organizational setting (e.g., employee performance review, hiring). Their work builds upon the theoretical framework of organizational justice by Colquitt (2001), which includes four dimensions: procedural fairness, interpersonal fairness, informational fairness (i.e., explainability), and distributive (i.e., outcome-based) fairness. However, only two of their four fairness dimensions (distributive fairness, and to a lesser extent, procedural fairness) are frequently discussed within the fairness literature of the AI community, as comprehensively reviewed by Pessach and Shmueli (2022). Furthermore, Ryan, Nadal, and Doherty (2023) interviewed roughly equal numbers of AI and HCI experts using procedural fairness and (four mathematical variations of) distributive fairness. They found an interesting theme among the HCI experts’ (but not the AI experts) responses: “Meeting a mathematical definition of fairness does not mean the model is ethical”. This finding suggests that, when humans interact with an AI prediction system to make decisions, i.e., HCI is involved, fairness research should extend beyond technical formulation to consider the social and legal background where the system is deployed.
HCI/AI vs. Law
Legal and technical AI/HCI research on conceptualizations of fairness have evolved mostly on their own, evidenced in the communities having studied similar or neighboring phenomena yet invented different terminologies and frameworks. For example, a well-received U.S. law review article on RRA fairness maps “disparate impact” to outcome fairness, and “disparate treatment” both to process (i.e., procedural) fairness and loosely to individual fairness (Mayson 2019). This mapping implies that individual fairness is not a subcategory of outcome fairness and that individual fairness is more related to procedural fairness than to group fairness. Both are inconsistent with the technical literature. Another seminal work is the Legal chapter (“Understanding United States anti-discrimination law”) in the Fair Machine Learning book by Barocas, Hardt, and Narayanan (2023). This book chapter, however, focuses its investigation of fairness concepts from the perspective of law review articles, and not primary sources of U.S. legal authority (e.g., case law). This is a methodology gap we seek to fill. While Yang and Dobbie (2020) combined thorough law review and statistical approaches, their work interprets the “Equal Protection” clause mostly with procedural fairness and may ignore outcome fairness criteria.
Our work contributes to the growing literature on reconciling law and algorithmic fairness. For example, Grossman, Nyarko, and Goel (2024) evaluated different statistical approaches to assess disparate impact. In addition, Xiang and Raji (2019) demonstrated examples of misalignment between ML fairness concepts and legal definitions.
Methodology – Legal Research
In order for technical AI fairness concepts to be adopted into actual RRA audit policies and survive judicial review, we conduct legal research to map these concepts to constitutional concepts.
Legal research can be analogized as a use case of the “document analysis” method by Bowen (2009), which is widely recognized in HCI. Document analysis includes four main steps: 1. “finding” documents, 2. “selecting” documents, 3. “appraising” (making sense of) documents, and 4. “synthesizing” data contained in documents. For legal research, “documents” include primary sources of U.S. federal and state law, e.g., the U.S. Constitution, case law (judicial branch), statutes (legislative branch), or regulations (executive branch) (Barkan, Mersky, and Dunn 2009). Linos and Carlson (2017) laid out common legal research steps (case sampling, case selection, case analysis, and developing theoretical explanation of the case outcomes), which are analogous to the four main steps (documents “finding”, “selecting”, “appraising”, and “synthesizing”) identified by Bowen (2009). We apply these four steps to identify relevant court cases and other primary sources of law.
Legal Research in HCI
Legal research, or legal document analysis, was employed in several HCI studies to inform design practices. Comber and Rossitto (2023) analyzed proceedings of a legal dispute in Sweden to study government vs. individual responsibility tension over waste management due to new technology, thereby informing the design of environmental responsibility frameworks among different stakeholders. Gray et al. (2021) synthesized several sources of European laws (the GDPR, the ePrivacy Directive, and the Court of Justice of the EU) to define what constitutes a “valid consent”, which serves as a key concept for their subsequent investigation on how “dark patterns” (i.e., implicit design to nudge users into selecting a certain privacy setting) in online consent banners should be regulated. Delgado (2020) analyzed judicial commentary (e.g., court opinions and law review articles) to understand how Technology-Assisted Review (TAR) during the discovery phase was perceived in the U.S. civil litigation community. Singh and Jackson (2021) adopted the document analysis method by Bowen (2009) on Indian legal documents to study how the Indian government developed Aadhaar, a biometrics-based identification project, to standardize their citizens’ data and deliver government services.
How?
For the first step (“finding” documents), we used a popular U.S. legal research database and search engine, Westlaw, to find primary legal sources with keyword search (such as “fairness”, “Equal Protection”, or “AI”), filters (case law, statutes, regulations), and legal citations in each case found from Westlaw to find more relevant past cases and observe how U.S. courts’ positions on the same legal question evolves over time. For the second step (“selecting” documents), we read the Westlaw-provided summary for each case law or skim through the first few paragraphs of each constitutional provision/statute/regulation to decide if they are actually relevant for our research questions. To ensure that the documents we selected still have legal effects, we also removed any sources of law that have been overturned by later appellate court decisions (marked with a red flag on Westlaw). For the third step (“appraising” documents), we wrote a summary for each selected legal document with relevant excerpts, clarified its legal authority (e.g., binding over which jurisdictions), and substantiated how it is related to fairness, recidivism risk assessment, or AI-assisted decision-making in general (deliverables from this step are in the Appendix). For the final step (“synthesizing data”), we discussed our theoretical findings in the next four sections.
Interaction among U.S. sources of law
As explained by Phillips (2019), in the U.S., there are four main sources of law at the federal level: the U.S. Constitution (the most binding authority), statutes (from the legislative branch), regulations (from the executive branch) which are more detailed implementation of statutes, and case law (from the judicial branch which follows a hierarchy in order of increasing bindingness: district courts, circuit courts of appeal, and the court of last resort or the Supreme Court) which interpret the Constitution and determine if any statutes or regulations are unconstitutional. Regarding case law, one important nuance is that a decision from a higher court is not binding for a lower court unless the lower court is within the (often geography-defined) jurisdiction of the higher court. For example, case law by the Second Circuit Court of Appeals are only binding for federal district courts within three states (Connecticut, New York, Vermont) and not for federal district courts located in other states. An analogous system of legal sources can be found at the state level in most U.S. states. The main principle of interaction between federal and state laws is that the rights that federal laws offer to individuals is a lower bound for state laws, i.e., states can grant their people more, but not fewer rights. Criminal law is a legal area reserved mostly for state laws, except for matters that cross the state borders, e.g., internet and airplane crimes, which fall under the federal jurisdiction (Samaha 2016). Therefore, in the context of RRA, there exist two parallel sets of state and federal laws.
Procedural Fairness and Outcome (Group and Individual) Fairness Are All Constitutionally Relevant
At a high level, the U.S. Constitution promotes fairness in criminal justice and other legal areas via two concepts: “Due Process” and “Equal Protection”. Relevant excerpts are “No person shall be […] deprived of life, liberty, or property, without Due Process of law” (the Fifth Amendment) and “…nor shall any State deprive any person of life, liberty, or property, without Due Process of law; nor deny to any person within its jurisdiction the Equal Protection of the laws” (the Fourteenth Amendment). Due process can be mapped to procedural (or process) fairness, e.g., which features should be used in an AI prediction model. Due to its straightforwardness, procedural fairness has been recognized by several low-level legal sources like statutes and regulations. As an example statute from the legislative branch, Section 3142(g) in the Bail Reform Act of 1984 specifies a long list of “factors to be considered” by judicial officers when making bail decisions, e.g., “the history and characteristics of the person” like employment and criminal history (Bail Reform Act 1984). From the executive branch, several state regulations explicitly enumerate allowed “risk factors”, e.g., age, sex/gender, number of prior convictions (Pennsylvania Administrative Code 2019a; California Code of Regulations 2010), or arrests/charges/convictions that mandate sex-offender registration (Oregon Administrative Code 2015b).
From the technical side, Grgić-Hlača et al. (2018) seminally define procedural (process) fairness as “the fraction of all users who consider the use of every feature in $F$ [the set of input features] to be fair” and mathematically developed three procedural fairness metrics based on the aforementioned definition and further conditioning: “feature-apriori fairness” (users having no prior knowledge of how the feature usage affects outcomes), “feature-accuracy fairness” (users knowing that using the feature increases accuracy), and “feature-disparity fairness” (users knowing that using the feature increases outcome disparity). Our conceptualization of procedural fairness (whether a feature should be included in the input space of RRA tools) aligns with “feature-apriori fairness” by Grgić-Hlača et al. (2018).
However, procedural fairness compliance is difficult to evaluate, especially in the context of recidivism risk assessment tools as many models are proprietary, some of which are proprietary and only subject to experts’ investigation of the algorithm if ordered by a judge, within the timeframe of a trial, and bound by several security protocols set by the tool owners, e.g., strict confidentiality constraints imposed by COMPAS owners on the expert witness Dr. Rudin in Flores v. Stanford (2021) (District Court, SD New York 2021). In the recent RRA landscape, the majority of tools used by state governments come from private companies. [^2] Legal Tech News (2020). The Most Widely Used Risk Assessment Tool in Each U.S. State. https://www.law.com/legaltechnews/2020/07/13/the-most-widely-used-risk-assessment-tool-in-each-u-s-state/. Therefore, in the long run, courts may help facilitate procedural fairness standards in the RRA context by clarifying to what extent RRA tool deployers may invoke trade secret standards in criminal proceedings. For example, Wexler (2018) argues that applying trade secret-based evidentiary privilege (to not disclose the technical inner-workings of RRA tools to the other parties during discovery), which were developed for the civil law context, directly to the criminal law context is “both harmful and unnecessary.” Furthermore, many theoretical rationales behind trade secret laws (e.g., to “facillitate controlled information sharing” with the government for public goods, and to incentivize owners of trade secrets to potentially turn them into disclosed patents) may in fact encourage trade secret disclosure to RRA decision subjects (defendants) through “narrow criminal discovery and subpoena powers combined with protective orders.”
However, it seems more practical, at least in the short run, to evaluate outcome fairness, which might correspond to the “Equal Protection” clause in the Fourteenth Amendment because this clause is the cornerstone of many discussions against group-based classification and subordination (Balkin and Siegel 2003). As a reminder, within outcome fairness, the AI Fairness literature covers two main schools of fairness: group fairness and individual fairness.
In U.S. law, case law interprets the “Equal Protection” clause as covering both group fairness and individual fairness. Regarding group fairness, the U.S. Supreme Court rules in Washington v. Davis (1976) that group fairness (“disproportionate impact”) matters but there should be another relevant school of fairness: “We have not held that a law […] is invalid under the Equal Protection Clause simply because it may affect a greater proportion of one race than of another. Disproportionate impact is not irrelevant, but it is not the sole touchstone of an invidious racial discrimination forbidden by the Constitution” (U.S. Supreme Court 1976b). As group fairness is ‘not the sole’ criterion here, we trace back to older U.S. Supreme Court decisions to find the missing piece. For instance, F.S. Royster Guano Co. v. Commonwealth of Virginia (1920) interprets “Equal Protection” consistently with individual fairness if we define “similarly situated” as having a high similarity function score: “The Equal Protection Clause of the Fourteenth Amendment commands […] essentially a direction that all persons similarly situated should be treated alike” (U.S. Supreme Court 1920).
Royster is in the civil law context, so we examine U.S. Supreme Court’s criminal law cases that also use the “similarly situated [individuals]” language when interpreting the constitutional “Equal Protection” concept: “Under the Equal Protection component of the Fifth Amendment’s Due Process Clause, […] to establish a discriminatory effect in a race case, the claimant must show that similarly situated individuals of a different race were not prosecuted” (U.S. v. Armstrong, 1996 (U.S. Supreme Court 1996)); “The imposition of the death sentence upon petitioner pursuant to the new statute did not deny him Equal Protection of the laws. Having been neither tried nor sentenced prior to Furman, he was not similarly situated to those whose death sentences were commuted” (Dobbert v. Florida, 1977 (U.S. Supreme Court 1977)). This line of individual fairness interpretation for the ‘Equal Protection’ clause is reaffirmed by more U.S. Supreme Court decisions such as Plyler v. Doe (1982).
Interestingly, a synergy of the relevance of the “Equal Protection” clause to both group fairness and individual fairness is provided in City of Cleburne, Tex. v. Cleburne Living Center (1985) where the U.S. Supreme Court rules that “Discrimination, in the Fourteenth Amendment sense, connotes a substantive constitutional judgment that two individuals or groups are entitled to be treated equally with respect to something” (U.S. Supreme Court 1985). Therefore, the “Equal Protection” clause in the U.S. Constitution has been interpreted as promoting both outcome fairness notions (group and individual fairness).
Specific Policies for RRA Touch On Procedural and Group Fairness, but not Individual Fairness
Athough some states such as California have developed elaborate legislation to control the quality of state-wide AI prediction tools with respect to group fairness metrics, we find limited evidence for individual fairness enforcement. Take for example Section 1320.35 of the California Penal Code, which from our search of AI-based pretrial risk assessment sources of law is the most fairness-detailed piece of legislation among U.S. states. However, most low-level fairness-related requirements in this statute are about group fairness only, requiring information about “risk levels aggregated by race or ethnicity, gender, offense type, ZIP Code of residency, and release or detention decision”, “the predictive accuracy of the tool by gender, race or ethnicity, and offense type”, and “any disparate effect in the tools based on income level”(California Penal Code 2021). The evaluation of metrics such as risk levels (output rate), predictive accuracy, and impact across sensitive features (race or ethnicity, sex/gender) and proxy features (those that highly correlate with sensitive features) shows that California cares about mitigating not only direct bias but also indirect bias. For example, even if an AI model does not use the sensitive feature race, the real-world bias can still penetrate into the model’s prediction via the use of features that highly correlate with race, e.g. ZIP Code of residence (Corbett-Davies et al. 2023). Interestingly, an AI explainability angle is also present here as the statute requires validation information about “line items, scoring, and weighting, as well as details on how each line item is scored”, or how much weight a prediction attributes to each feature. This explainability requirement exposes the inner decision-making process of the model, so it can be linked to procedural fairness. However, individual fairness is not relevant in this statute. Another example is Section 725 ILCS 5/110-6.4 of the Illinois Compiled Statutes. It requires their state-wide risk assessment tool to “not discriminate on the basis of race, gender, educational level, socio-economic status, or neighborhood”, which can only be attributed to group fairness, not individual fairness (Illinois Compiled Statutes 2023).
Procedural fairness considerations can also be found in regulations (executive branch). For example, 204 Pa. Code § 305.2 (Pennsylvania Administrative Code 2019b) explicitly supports using age and sex/gender as RRA model inputs.
In summary, while the three technical fairness concepts (procedural, group, and individual fairness) may have a constitutional basis, among them, only procedural fairness and group fairness can be found in specific statutes and regulations, as shown in Figure 1.
Figure 1: Whether Individual, Group, and Procedural Fairness Concepts are recognized by the U.S. Constitution, case law (judicial), statutes (legislative), or regulations (executive)
Demographic Features in the Same Judicial “Scrutiny” Category Might Trigger Slightly Different Scrutiny by U.S. Courts in Practice
As procedural fairness is found in concrete statutes and regulations, it is not surprising that the relationship between demographic features and procedural fairness has been well established in case law. For example, in Clark v. Jeter (1988), the U.S. Supreme Court acknowledges a common law hierarchy of features (classes or group memberships) on which the government may discriminate to varying degrees for the sake of public interest, from the most to the least stringent amount of justification the government must provide: “strict scrutiny” (e.g., race and alienage/national origin), “intermediate scrutiny” or “heightened scrutiny” (e.g., sex/gender and legitimacy/out-of-wedlock status), and “rational basis” (e.g., age and disability) (U.S. Supreme Court 1988).
In particular, at one end of the judicial scrutiny-worthiness spectrum, a “strict scrutiny” feature means that a state or federal government’s policy that treats one group differently than another group based on that feature (e.g., race) must be the “least restrictive means” to further a “compelling government interest.” [^3] https://www.law.cornell.edu/wex/strict˙scrutiny At the other end of the spectrum, a ”rational basis” feature means that a government’s differential treatment policy based on that feature (e.g., age) must serve a “legitimate state interest”, and there must be a “rational” connection between the policy’s means and interest. [^4] https://www.law.cornell.edu/wex/rational˙basis˙test In the middle of the spectrum, an “intermediate/heightened scrutiny” feature means that government’s differential treatment policy based on that feature (e.g., sex/gender) must further an “important government interest” (a higher burden of proof than the “legitimate” interest in rational basis, but lower than the “compelling” interest in strict scrutiny), and the policy’s means must be “substantially related to” that interest (a higher burden of proof than a “rational connection”, but lower than the “least restrictive means”). [^5] https://www.law.cornell.edu/wex/intermediate˙scrutiny
Instead of viewing strict scrutiny, intermediate scrutiny, and rational basis as three discrete values of warranted scrutiny amount, we may instead formulate them as three continuous ranges because case law has shown that two features within the same group (or range) may not necessarily warrant the exact same amount of scrutiny.
For example, U.S. Supreme Court’s (SCOTUS) decisions recognize race and alienage as strict scrutiny features: 1/ Korematsu v. United States, 323 U.S. 214 (1944): “all legal restrictions which curtail the civil rights of a single racial group are immediately suspect. […] courts must subject them to the most rigid scrutiny.” (U.S. Supreme Court 1944). 2/ Graham v. Richardson, 403 U.S. 365 (1971): “classifications based on alienage, like those based on nationality or race, are inherently suspect and subject to close judicial scrutiny” (U.S. Supreme Court 1971). 3/ Bernal v. Fainter, 467 U.S. 216 (1984): “a state law that discriminates on the basis of alienage can be sustained only if it can withstand strict judicial scrutiny” (U.S. Supreme Court 1984).
However, case law shows that SCOTUS has scrutinized alienage-based discrimination with two slight de facto differences compared to race-based discrimination.
The first difference is that while SCOTUS mandates the same amount of (strict) scrutiny applies regardless of whether a race-based discrimination comes from a state or federal government, SCOTUS applies less scrutiny on alienage-based discrimination from the federal government than on alienage-based discrimination from state governments. In Hampton v. Mow Sun Wong (1976), SCOTUS hints that judicial scrutiny over alienage-based discrimination by the federal government in civil service hiring may be lower than scrutiny over alienage-based discrimination by a state government: “overriding national interests may justify a citizenship requirement in the federal service even though an identical requirement may not be enforced by a State” (U.S. Supreme Court 1976a). However, for race-based discrimination, in Adarand Constructors, Inc. v. Peña, 515 U.S. 200 (1995), SCOTUS reasons that “All racial classifications, imposed by whatever federal, state, or local governmental actor, must be analyzed by a reviewing court under strict scrutiny” (U.S. Supreme Court 1995).
The second difference is a “political function” exception, which enables state/federal government to justify alienage-based discrimination but not race-based discrimination. In Bernal v. Fainter, 467 U.S. 216 (1984), “We [SCOTUS] have, however, developed a narrow exception to the rule that discrimination based on alienage triggers strict scrutiny […] and applies to laws that exclude aliens from positions intimately related to the process of democratic self-government” (U.S. Supreme Court 1984). SCOTUS used this “political function” exception to avoid striking down state policies that exclude aliens (non-U.S. citizens) from public service jobs such as police officers in Foley v. Connelie (U.S. Supreme Court 1978), teachers in Ambach v. Norwick (U.S. Supreme Court 1979), and probation officers in Cabell v. Chavez-Salido (U.S. Supreme Court 1982).
In conclusion, case law supports our decision to model each scrutiny standard (e.g., strict scrutiny) as a continuous-valued range (instead of a discrete category) to accommodate several demographic features that may slightly differ in de facto scrutiny-worthiness (e.g., alienage and race).
Jurisdictions Disagree on Procedural Fairness Criteria (Which Demographics to Use as Model Inputs)
Although the scrutiny-based hierarchy is not legally binding for the recidivism context, there is consensus not to use features in the “strict scrutiny” range as model inputs. The consensus on race is clear. Although most risk assessment tools, e.g., COMPAS, do not use race as a feature, many works such as Johnson (2021) go further to criticize the tool for indirectly perpetuating racial bias through proxy features (those highly correlated with race). Hu (2023) gives a constructivist argument for why acting on proxy features of race is not different than acting on the basis of race. Regarding alienage/national origin, a sample questionnaire [^6] https://www.documentcloud.org/documents/2702103-Sample-Risk-Assessment-COMPAS-CORE shows that COMPAS does not collect this feature, and we do not find any literature discussing the use of or unfairness with respect to this feature by any recidivism tools, indicating an implicit consensus that national origin should not be used.
For the lower two ranges (intermediate scrutiny and rational basis), there remains disagreement on whether features at those ranges might be used. The first example of disagreement is sex/gender (at the intermediate scrutiny range). On the one hand, the Wisconsin Supreme Court in their State v. Loomis (2016) decision strongly advocates the use of sex/gender as a feature: “COMPAS’s use of gender promotes accuracy that ultimately inures to the benefit of the justice system, including defendants” (Wisconsin Supreme Court 2016). On the other hand, Section 2A:162-25(2) of the New Jersey Statutes adopts a clear stance against using sex/gender as a feature for AI recidivism risk assessment: “Recommendations for pretrial release shall not be discriminatory based on race, ethnicity, gender, or socio-economic status” (New Jersey Statutes 2024).
The second example of disagreement is age (at the rational basis range). For example, in the same jurisdiction of New York, while Section 168-l of the Consolidated Laws of New York explicitly requires sex offense recidivism risk assessment to take into account “the age of the sex offender at the time of the commission of the first sex offense” (Consolidated Laws of New York 2011), in their Flores v. Stanford (2021) decision, the U.S. district court for Southern District of New York indicates an implicit stance against the use of age in recidivism risk assessment by allowing expert inspection of the data used to train COMPAS to “help Plaintiffs substantiate their allegations that COMPAS punishes juvenile offenders for their youth, such that Defendants’ reliance on this tool is constitutionally problematic” (District Court, SD New York 2021).
Not only case law and statutes, but state regulations also disagree. While some regulations include both sex/gender and age as risk factors (e.g., 204 Pa. Stat. Ann. § 305.2 (Pennsylvania Administrative Code 2019b) and Cal. Code Regs. tit. 15, § 3768.1 (California Code of Regulations 2010)), other regulations include neither sex/gender nor age (e.g., 20 Ill. Adm. Code 1905.60 (Illinois Administrative Code 2017) and Or. Admin. R. 291-078-0020 (Oregon Administrative Code 2015a)).
In summary, while there is a consensus that strict scrutiny features should be excluded from model inputs, legal sources disagree on intermediate scrutiny (sex/gender) and rational basis (age) features.
Proposal: Integrating Fairness Criteria and Demographic Features into a Scrutiny-based Framework
The controversy on whether features in the two lower ranges of scrutiny can be used as input for recidivism risk assessment models indicates room for further research that links to the procedural vs. outcome fairness and the individual versus group fairness dichotomies. In the same framework of three continuous scrutiny ranges, we propose three discrete scrutiny thresholds: First is the procedural fairness or “exclusion from model inputs” threshold (if we take the view by Lee et al. (2019); Agan and Starr (2018) that excluding a feature from the model input space corresponds to procedural fairness). Second is the group fairness or “group parity required” threshold. Third is the individual fairness or “ignorance in individual similarity function” threshold. The more descriptive (and lengthier) name of each threshold corresponds to a pass condition, i.e., a feature’s scrutiny amount warranted is higher than this threshold if people believe that the feature should satisfy the corresponding threshold’s pass condition.
First, a feature that people believe should be excluded from the input space of an RRA model will have its scrutiny score passing (above) the procedural fairness threshold. Second, for a feature, if people believe it should be the reference feature so that they can compute a group parity metric and require that parity metric to be near perfect, then this feature’s scrutiny score will pass the group fairness threshold. Finally, if people believe that to evaluate an individual fairness metric, the individual similarity function must ignore a certain feature, then this feature’s scrutiny score will pass the individual fairness threshold.
It remains for future work to investigate whether it is normatively plausible to, and if so, how to rank these three fairness thresholds against one another. A potential, unvalidated ordering of them is illustrated in Figure 2. Note that satisfying a higher threshold does not mathematically guarantee satisfying a lower threshold, e.g., perfect outcome parity across races does not guarantee that race is excluded from model inputs. Rather, any potential ordering is only a normative judgment on which threshold should reflect higher judicial scrutiny.
We provide some justifications for this not yet empirically validated ordering: An example of why the group fairness threshold (orange) is above the procedural fairness threshold (yellow) is the recurrent debate on whether COMPAS should achieve racial parity (i.e., race may pass or fail the orange threshold) even though COMPAS does not use race as a predictive feature (i.e., race passes the yellow threshold) (Washington 2018). An intuition for why the individual fairness threshold (red) is above the group fairness threshold (orange) is a thought experiment: even if one believes that a model should not have disparate racial impact (i.e., race passes the orange threshold), one might still believe that all else equal, two same-race individuals are more similar than two different-races individuals (i.e., race fails the red threshold). Thinking reversely, suppose one believes that race is not a good proxy to measure criminal/violence prevalence (or equivalently, all races have the same amount of criminal/violence prevalence), and therefore, race should be excluded from the pairwise individual similarity function (i.e., race passes the red). This premise should imply that the recidivism risk should be independent of race or equalized across racial groups (i.e., race passes the orange threshold). In other words, passing the red threshold might imply normatively, but not mathematically, passing the orange threshold, indicating that the individual fairness threshold reflects higher scrutiny than the group fairness threshold. However, this ordering has not been empirically validated.
Contextualizing our fairness thresholds in the judicial scrutiny-based framework, we only know that the procedural fairness threshold must be below the strict scrutiny range because all features in this range (race, national origin) are excluded from model inputs. As legal sources disagree on whether sex/gender and age should be excluded from model inputs, we do not know where our procedural fairness threshold should go with respect to the intermediate scrutiny and rational basis ranges. We illustrate this open question in Figure 3.
Figure 2: Potential ordering of proposed scrutiny thresholds.
Figure 3: Legal sources classify demographic features into three scrutiny ranges. For RRA, only strict scrutiny features are above the “exclusion from model inputs” (yellow) scrutiny threshold, i.e., expected by legal sources to be excluded from the AI models’ input space. There are conflicting legal sources about whether intermediate scrutiny or rational basis features (such as sex/gender or age) should be excluded.
Figure 4: How the empirical findings by Nguyen et al. (2025) might be applied to our scrutiny-based framework.
Suppose empirical works, e.g., human studies with laypeople or court officials, can confirm that our proposed three scrutiny thresholds are consistent with humans’ fairness perception, our framework might help systematically determine whether procedural, group, and/or individual fairness criteria should be audited with respect to each demographic feature.
One empirical work highly relevant to our framework is the human subjects experiment conducted by Nguyen et al. (2025), which tests whether including a demographic feature (race, sex/gender, or age) as available information for laypeople to evaluate pairwise similarity between criminal justice defendants would worsen those laypeople’s individual fairness judgement (i.e., making them more likely to judge opposite risk scores on similar defendants as fair). They conclude that the pairwise individual similarity function should include sex and age but ignore race. They also have a short survey question on procedural fairness and find that most laypeople prefer demographic features (race, sex/gender, and age) to be excluded from the RRA model inputs.
Figure 4 summarizes how the human subjects experiment by Nguyen et al. (2025) might further clarify our proposed framework. Regarding the “exclusion from model inputs” (yellow) threshold, most of their participants believe criminal history features should be included (i.e., below the yellow threshold) and a majority of participants think that age, sex/gender, and race should be excluded (i.e., above the yellow threshold). Regarding the “ignorance in individual similarity function” (red) threshold, their empirical results and analyses conclude that the pairwise individual similarity function can consider age and sex/gender (i.e., below the red threshold), but it should ignore race (i.e., above the red threshold). Note that due to the relative positions of the red and yellow thresholds compared to sex/gender and age (red >>> { sex/gender, age } >>> yellow), applying the transitive property of inequality, the red threshold is above the yellow threshold, partially validating our theoretically proposed ordering. Finally, the three (blue) scrutiny ranges from legal precedents (strict scrutiny, intermediate scrutiny, rational basis) are placed around each respective demographic feature. However, their results do not confirm whether the red threshold falls within the strict or intermediate scrutiny range.
As a caution, this is just one out of many possible empirical arrangements of our proposed framework. For example, another possible arrangement is that sex/gender and age be placed below the yellow (procedural fairness) threshold because Katsiyannis et al. (2018) found sex/gender and age to be predictive of recidivism. Therefore, we invite future empirical works to refine our judicial scrutiny-based framework of fairness criteria and demographic features.
International Perspectives: How our U.S. Law Findings Relate to Other Major Jurisdictions
We contextualize our U.S. legal research findings and proposed scrutiny-based framework in other jurisdictions where there is a strong AI market and evidenced/potential use of AI-assisted decision-making tools by government bodies.
The European Union
The EU AI Act categorizes RRA tools (”Individual criminal offence risk assessment or prediction”) as “unacceptable risk” and therefore bans it. However, according to Recital 42: “risk assessments carried out with regard to natural persons in order to assess the likelihood of their offending or to predict the occurrence of an actual or potential criminal offence based solely on profiling them or on assessing their personality traits and characteristics should be prohibited. In any case, that prohibition does not refer to or touch upon risk analytics that are not based on the profiling of individuals or on the personality traits and characteristics of individuals, such as AI systems using risk analytics to assess the likelihood of financial fraud by undertakings on the basis of suspicious transactions” (The European Parliament and Council 2024). This likely means that any AI-assisted decision making tools used by EU member states are prohibited from including any demographic features (e.g., race, sex, or age) as model input features, i.e., all demographic features are placed above the “exclusion from model inputs” threshold if we adapt our scrutiny-based framework to the EU context.
Prior to the EU AI Act, as reviewed by Weerts et al. (2023), four main directives lay out the theoretical framework for EU non-discrimination law, which also applies to algorithmic fairness considerations: the Race Equality Directive 2000/43/EC (Council of European Union 2000b), the Framework Equality Directive (Council of European Union 2000a), the Gender Equality Directives 2004/113/EC (Council of European Union 2004) and 2006/54/EC (Council of European Union 2006). This EU framework defines two types of fairness concerns: direct discrimination (”one person is treated less favourably than another is, has been or would be treated in a comparable situation on grounds of” a protected feature like race, sex/gender, and age) and indirect discrimination (“an apparently neutral provision, criterion or practice would put persons of a protected group at a particular disadvantage”). While the strictly forbidden “direct discrimination” directly reaffirms our procedural fairness observation from the EU AI Act that race, sex, or age should be excluded from model input features, the less restricted “indirect discrimination” under EU non-discrimination law hints at our “group parity required” scrutiny threshold. Interestingly, “indirect discrimination” might be legal if it “is objectively justified by a legitimate aim and the means of achieving that aim are appropriate and necessary”, which was interpreted as “if the same legitimate aim can be achieved through less discriminatory alternatives, those must be used” (Weerts et al. 2023; Tobler 2005). This “[no] less discriminatory alternatives” language of permissible “indirect discrimnation” under EU non-discrimination law is highly similar to the “least restrictive means” language in the U.S. “strict scrutiny” review standard against discriminatory government policies. If an AI model with disparate impact (e.g., violating demographic parity against certain races) can survive “indirect discrimination” challenges under EU law, it is likely to survive the “strict scrutiny” test under U.S. law too.
Regarding individual fairness, Weerts et al. (2023) find the “likes should be treated alike” Aristotelian formulation of justice to be the philosophical foundation of the EU “direct discrimination” doctrine. However, after thoroughly reviewing specific and operationalizable AI-related sources of EU law such as the General Data Protection Regulation, GDPR (The European Parliament and Council 2016), the Digital Services Act, DSA (The European Parliament and Council 2022), and a proposal version of the EU AI Act, Calvi and Kotzinos (2023) conclude that “the few references to fairness metrics existing in the EU legal framework hint at group fairness and not individual fairness.” This problem is similar to our finding that although SCOTUS linked the U.S. constitutional “Equal Protection” clause to the technical individual fairness concept, we find no concrete individual fairness standards in U.S. federal/state statutes or regulations.
China
Although Chinese courts do not have an RRA use case, some of their criminal courts use AI to estimate the “social harm” of the current case based on similar past cases from other defendants (instead of predicting future criminal cases from the same defendant). Judges may consider AI-estimated “social harm” as “a frame of reference for sentencing,” which is analogous to U.S. judges considering RRA risk scores in sentencing (Papagianneas and Junius 2023; Cui 2020).
Regarding technical/legal fairness criteria, unlike the procedural vs. outcome fairness dichotomy in American/European laws, Papagianneas and Junius (2023) synthesize a procedural vs. substantive fairness dichotomy after reviewing judicial documents published by China’s Supreme People’s Court (SPC) to guide lower courts. Substantive fairness in China has nothing to do with outcome-based (group/procedural) fairness, but its goals are 1/ sustaining legitimacy (of the ruling party), 2/ maintaining social stability, and 3/ user convenience. On the other hand, procedural fairness in China is often less about a set of legally allowed/forbidden features for model inputs, but more about the explainability angle (internal accountability, procedural consistency, and external visibility), e.g., whether AI-assisted decision makers may explain to decision subjects what features they consider and why. Furthermore, the expected response to potentially discriminatory actions by a government agency in China is often to report to that agency or a higher government body for a top-down solution, rather than through bottom-up litigation against the government, i.e., judicial scrutiny of government actions is rare (Cui, Cheng, and Wiesner 2019). Therefore, our judicial scrutiny-based fairness framework is irrelevant to China’s legal system.
India
The most important supplement to our framework in India’s context is to expand the set of demographic features for fairness consideration to include “caste”, a hereditary Hindu class feature which historically determines one’s access to financial, cultural, social, and many other types of capital in life (Sambasivan et al. 2021). If placed on our legal framework, caste and sex/gender are likely above the “group parity required” threshold due to “reservations”, a constitutionally mandated practice to reserve certain quotas to access public resources like education and bank loans for vulnerable groups such as historically marginalized castes and women (Basavaraju 2009; Sambasivan et al. 2021). This policy seeks to equalize access to resources across demographic groups and is therefore related to group fairness. A recent case decided by the Supreme Court of India, LT. Col Nitisha v. Union of India (2021), though not in the AI context, further reaffirms group fairness protection for women as the Court recognizes indirect discrimination by the government (e.g., disparate impact of a policy on women, even without proven discriminatory intent) to be unconstitutional, just like direct discrimination (Supreme Court of India 2021).
Discussion
We contemplate how our legal findings and new framework might impact different stakeholders in the design and refinement of RRA and other AI-assisted decision-making tools.
Tool designers
Though not directly in the criminal justice context, some HCI scholars and practitioners started to consider fairness as part of the design pipeline. For example, Nakao et al. (2022) designed an interactive “explanatory debugging” interface to help non-expert end-users identify and potentially correct fairness issues in the loan application context. However, this is still a minority. For example, within the Mobile HCI community, Yfantidou et al. (2023) expressed concerns that only 5% of the Mobile HCI papers they reviewed conformed to fairness criteria, calling for more systematic fairness research in that sub-community. In HCI more broadly, some designers of AI-assisted decision-making tools like RRA might not have enough incentives to spend extra time and resources on auditing any potential bias in their products or mitigating discovered bias through algorithmic refinement or alternative data. Our legal research might provide designers with external, legally binding incentives to adopt fair and responsible AI practices, as Grimpe, Hartswood, and Jirotka (2014) showed that policy might facilitate the adoption of responsible design principles (e.g., inclusiveness) into design practices.
Some other designers might simply not have enough knowledge about the myriad of technical fairness criteria or fair design principles out there in the technical AI literature. By integrating technical fairness criteria and demographic features into a framework of scrutiny, we might streamline the fairness audit process to help designers understand what types of demographic data they need to collect and what types of fairness criteria they should optimize for.
Tool users
When a high-stakes AI tool makes bad (e.g., unfair) decisions, it is important to enable the tool users to recognize and correct such decisions. This problem of Algorithmic Recourse has gained increasing attention in HCI, especially Human-centered AI. For example, Esfahani et al. (2024) designed an interaction paradigm to elicit users’ preferences for algorithmic decisions and guide users towards effective recourse interventions for unfavorable decisions. In RRA and other high-stakes contexts, the users are often specialized experts who follow specific professional traditions, which algorithm recourse should also consider. For example, Yacoby et al. (2022) found that counterfactual explanations, a popular algorithmic recourse to uncover bias (Karimi, Schölkopf, and Valera 2021), were simply ignored by judges because judges perceived their job as deciding real cases for real people, not hypothetical ones.
Judges have little incentive to consider algorithmic recourse or avoid overreliance on RRA tools because, as long as a judge still makes the final bail/sentencing decision, it becomes a shield to protect their decision against accusations of fairness violation in future appeals. In Brooks v. Commonwealth, although there is evidence that the trial judge favors the AI risk assessment-based sentencing recommendation and dismisses the shorter active sentence in the non-AI recommendation, the Court of Appeals of Virginia rules that ‘the trial court properly exercised its discretion’ without inquiring the level of reliance by the trial court on the AI risk assessment (Court of Appeals of Virginia 2004).
Therefore, by attributing a strong legal basis to existing technical fairness criteria and our proposed scrutiny-based framework, we encourage future work on designing new recourse intervention methods based on these criteria and framework. Such legally grounded recourses might confront the judges with evidence of potential bias in the RRA tool more directly, incentivizing them to respond to such bias more promptly. For example, if a recourse can show that an RRA tool gave opposite risk scores to many pairs of defendants with different races (but similar criminal histories, sexes/genders, and ages), which we showed to violate the constitutionally relevant individual fairness criterion, judges will be more likely to suspend the use of that tool in their court rooms and require further fairness audits.
Tool decision subjects
Before one can find recourse for bad algorithmic decisions, one must be able to contest those. The contestability literature in HCI often assumes the one who should contest is the model user (Vaccaro et al. 2019). However, Lyons, Velloso, and Miller (2021) found that first, the primary goal of contestability is to protect individuals, and second, contestability in AI-assisted decision-making resembles contestability in human decision-making. Therefore, in RRA and other high-stakes contexts where the decision subjects (e.g., defendants) differ from the model users (e.g., judges), the primary focus of contestability design should be on the decision subjects because they have the most incentives to protect themselves and are separate from the decision makers, i.e., tool users (so that research on AI-assisted decision-making contestability can draw analogies from human decision-making contestability).
RRA decision subjects face a high imbalance of contestability power against RRA users because the burden of proof to show that judges over-rely on AI is often placed on the defendants. In People v. Younglove (2019), when the defendants contended that their right to an ‘individualized sentencing decision’ was violated by the COMPAS risk score presented to the sentencing judge, the Michigan Court of Appeals ruled that “defendants offer no evidence that their sentencing courts actually placed significant (or any) weight on the COMPAS assessments in crafting their sentences. Defendants have failed to carry their burden of showing that the inclusion of the information affected their substantial rights” (Michigan Court of Appeals 2019).
Defendants have almost no way to reconstruct the mental model of the judges’ decisions, especially when most judges are not required to disclose how much weight they give to the AI risk score, which may internalize demographics-based bias. For example, Stevenson and Doleac (2022) showed that, conditioned on the same recidivism risk level, AI-assisted judges in Virginia give black defendants sentences that are 15-20% longer than white defendants.
By integrating nuanced legal concepts of scrutiny and abstract fairness criteria into a simple framework, we might empower criminal defense attorneys to spot fairness problems of RRA tools used on their clients more effectively, thereby enhancing the clients’ contestability and potentially suing the tool designers for wrongful discrimination or personal liberty damages, which may incentivize fairer algorithmic design and refinement. An attorney is often the closest companion a defendant gets when going through the criminal justice system. Karusala et al. (2024) found that accompaniment, i.e., being accompanied by a professional who also cares about the decision subject, is critical for marginalized communities (those most susceptible to algorithmic bias) to contest AI decisions.
How Technical AI Fairness Works Further Inform Our Legal Framework
Rudin (2019) argues against using black-box RRA tools because even if one tries to interpret black-box models’ predictions with post hoc explanations (developing a second model to perturb the original model’s outcomes to explain it), those explanations are not necessarily faithful. Their advocating for inherently interpretable models in high-stakes contexts like RRA will enable easier audit of “procedural fairness” in our framework, as expert witnesses might investigate the source code to determine whether the model uses a demographic feature and with how much weight. Both works by Lagioia, Rovatti, and Sartor (2023) and Chouldechova (2017) demonstrate the potential incompatibility of several group fairness metrics (“a system that is equally accurate for different groups may fail to comply with group-parity standards”; “disparate impact can arise when an RPI fails to satisfy the criterion of error rate balance”), informing our future work to explore other variations of our group fairness-based (yellow) scrutiny threshold and their relative orders in our framework.
Conclusion
Our legal research shows that while major technical AI fairness criteria (procedural, group, and individual fairness) have constitutional bases, they have not yet been translated into consistent RRA statutes and regulations. To fill this gap, we extend the demographics-related scrutiny framework from U.S. case law to incorporate fairness criteria as scrutiny thresholds. Finally, we contextualize our framework in three other jurisdictions with a big AI market.
Appendix A Appendix: Detailed Review of Fairness-related Sources of U.S. Law
Before synthesizing the legal findings presented in this paper, we conducted a detailed review of the four main sources of federal and state laws: the U.S. Constitution, statutes, case laws, and regulations. In each category (source of law), we identify legal items that are either related to AI, recidivism risk assessment, or general fairness-related provisions (“equal protection” or “due process”), give a quick summary and analysis of each source, and quote the most relevant excerpt from the source, as presented in the 30-pages attachment below. Please note that not all legal items below go into the main narrative of the paper.